# Efficient Uncertainty Propagation in Bayesian Two-Step Procedures

### v1

@@ 4 Conclusion
In this work, we introduced a principled and computationally efficient method for uncertainty propagation in Bayesian two-step modeling frameworks. Our approach combines Pareto Smoothed Importance Sampling (PSIS) and Importance Weighted Moment Matching (IWMM) to significantly reduce the number of costly sampling runs typically required for reliable uncertainty propagation. Rather than performing separate MCMC simulations for each of the multiple realizations of the uncertainty quantity, we run MCMC for only a small, carefully selected subset of representative draws and approximate the remaining posteriors via the two importance sampling techniques. We focused on two important application domains where uncertainty propagation is particularly critical: multiple imputation and surrogate modeling.
@@
Despite its strong performance across a wide range of scenarios, the method shows limitations in particularly challenging settings. For example, in missing data problems characterized by complex posterior structure – such as those arising from shrinkage priors, high-dimensional parameter spaces, and high proportion of missingness – the affine transformations used in IWMM are insufficient. In such cases, the discrepancy between the target and proposal distribution leads to degraded approximation quality. More expressive transformation techniques may be required, similar to these needed in other applications of IWMM, such as leave-one-group-out cross-validation [69].
Although we demonstrated the method in two specific domains, its applicability extends to any modeling scenario with a two-step structure where uncertainty propagation is essential. Developing scalable and user-friendly implementations will be crucial for broader adoption. While the current implementation is mostly written in R (see GitHub), versions in a compiled language such as C++, or full integration with probabilistic programming frameworks like Stan could greatly enhance practical performance.
@@ Acknowledgements
This research was supported by the German Research Foundation (DFG) via Collaborative Research Cluster 391 "Spatio-Temporal Statistics for the Transition of Energy and Transport" – 520388526.
@@ Appendix B Iterative Algorithm for using a Mixture Proposal
Input: Variable of interest $\theta\sim p(\theta)$, underlying variable $\tau\sim p(\tau)$, mixture size $I_{mix}$
1 Draw $m$ samples $\tau^{(1)},\dots,\tau^{(m)}\sim p(\tau)$.
2 Set index set $I=\{1,\dots,m\}$.
3 while $|I|>|I_{mix|}$ do
4       Select $|I_{mix}|$ representative values from $\{\tau^{(i)}\}_{i\in I}$ and run MCMC to access their posterior distributions through samples $\{\theta^{(s)}_{i}\}_{s=1}^{S}$. Build the mixture distribution $p_{\rm mix}(\theta\mid\{\tau^{(i)}\})=\frac{1}{|I_{mix}|}\sum_{i\in I_{mix}}p(\theta|\tau^{(i)})$ and get draws $\{\theta^{(s)}_{*}\}_{s=1}^{S}\sim p_{\rm mix}(\theta\mid\{\tau^{(i)}\})$ trough resampling.
5       Remove the $|I_{mix}|$ selected values from the index set $I$.
6       for $i\in I$ do
7             Run PSIS with target $p(\theta\mid\tau^{(i)})$ and proposal $p_{\rm mix}(\theta\mid\{\tau^{(i)}\})$ and gain importance weights and metric $\hat{k}_{i}$.
8             if $\hat{k}_{i}<0.7$ then
9                   Use calculated importance weights and draws $\{\theta^{(s)}_{*}\}_{s=1}^{S}\sim p_{\rm mix}(\theta\mid\{\tau^{(i)}\})$ for importance resampling and save the resampled draws as posterior draws corresponding to $p(\theta\mid\tau^{(i)})$.
10                   Remove $i$ from $I$: $I\leftarrow I\setminus i$.
11
12
13
14for $i\in I$ do
15       Run MCMC to gain posterior draws from $p(\theta\mid\tau^{(i)})$.
16
17Calculate the marginalized posterior via Monte Carlo Estimate $p(\theta\mid C)\approx\frac{1}{m}\sum_{i=1}^{m}p\left(\theta\mid\tau^{(i)}\right)$.
18
Algorithm 2 Iterative Algorithm for efficiently calculating the marginalized posterior via the Monte Carlo Estimate (mixture proposal)

### v2

@@ 4 Conclusion
In this work, we introduced a principled and computationally efficient method for uncertainty propagation in Bayesian two-step modeling frameworks. Our approach combines Pareto Smoothed Importance Sampling (PSIS) and Importance Weighted Moment Matching (IWMM) to significantly reduce the number of costly sampling runs typically required for reliable uncertainty propagation. Rather than performing separate MCMC simulations for each of the multiple realizations of the uncertainty quantity, we run MCMC for only a small, carefully selected subset of representative draws and approximate the remaining posteriors via the two importance sampling techniques. We focus on two important application domains where uncertainty propagation is particularly critical: multiple imputation and surrogate modeling.
@@
Despite its strong performance across a wide range of scenarios, the method shows limitations in particularly challenging settings. For example, in missing data problems characterized by complex posterior structure — such as those arising from shrinkage priors, high-dimensional parameter spaces, and high proportion of missingness — the affine transformations used in IWMM are insufficient. In such cases, the discrepancy between the target and proposal distribution leads to degraded approximation quality. More expressive transformation techniques may be required, similar to these needed in other applications of IWMM, such as leave-one-group-out cross-validation [69].
Although we demonstrated the method in two specific domains, its applicability extends to any modeling scenario with a two-step structure where uncertainty propagation is essential. Bayesian cut models are one example of such a setting. Developing scalable and user-friendly implementations will be crucial for broader adoption. While the current implementation is mostly written in R (see GitHub), versions in a compiled language such as C++, or full integration with probabilistic programming frameworks like Stan could greatly enhance practical performance.
Choosing $m$, the number of first-step draws $\tau^{(i)}$, is inherently context-dependent and often difficult in practice. Even in the closely related importance-sampling literature, the choice of $m$ has been debated for decades, underscoring that there is no universally optimal rule. Our approach alleviates this trade-off by making it feasible to use larger values of $m$ and thus propagate first-step uncertainty more faithfully, without incurring a proportional increase in computational cost.
@@ Acknowledgements
This research was supported by the German Research Foundation (DFG) via Collaborative Research Cluster 391 ”Spatio-Temporal Statistics for the Transition of Energy and Transport” – 520388526.
@@ Appendix B Iterative Algorithm for using a mixture proposal
Algorithm 2 Iterative Algorithm for Efficiently Calculating the Marginalized Posterior (Mixture Proposal)
1:Variable of interest $\theta\sim p(\theta)$, underlying variable $\tau\sim p(\tau)$, mixture size $I_{\text{mix}}$
2:Draw $m$ samples $\tau^{(1)},\dots,\tau^{(m)}\sim p(\tau)$
3:Set index set $I=\{1,\dots,m\}$
4:while $|I|>|I_{\text{mix}}|$ do
5:  Select $|I_{\text{mix}}|$ representative values from $\{\tau^{(i)}\}_{i\in I}$
6:  Run MCMC to access their posterior distributions: $\{\theta^{(s)}_{i}\}_{s=1}^{S}$
7:  Build the mixture:
$p_{\text{mix}}(\theta\mid\{\tau^{(i)}\})=\frac{1}{|I_{\text{mix}}|}\sum_{i\in I_{\text{mix}}}p(\theta\mid\tau^{(i)})$
8:  Obtain samples $\{\theta^{(s)}_{*}\}_{s=1}^{S}\sim p_{\text{mix}}(\theta\mid\{\tau^{(i)}\})$ via resampling
9:  Remove the $|I_{\text{mix}}|$ selected values from II
10:  for $i\in I$ do
11:   Run PSIS with target $p(\theta\mid\tau^{(i)})$ and proposal $p_{\text{mix}}(\theta\mid\{\tau^{(i)}\})$
12:   Obtain importance weights and diagnostic metric $\hat{k}_{i}$
13:   if $\hat{k}_{i}<0.7$ then
14:     Use the importance weights and $\{\theta^{(s)}_{*}\}$ to resample posterior draws for $p(\theta\mid\tau^{(i)})$
15:     Remove $i$ from II: $I\leftarrow I\setminus\{i\}$
16:   end if
17:  end for
18:end while
19:for $i\in I$ do
20:  Run MCMC to obtain posterior draws from $p(\theta\mid\tau^{(i)})$
21:end for
22:Compute marginalized posterior via Monte Carlo estimate: $p(\theta\mid C)\approx\frac{1}{m}\sum_{i=1}^{m}p(\theta\mid\tau^{(i)})$
@@ Appendix C Additional Results of the Simulation Study for the Multiple Imputation Problem
Figure 12: Posterior density estimates (MCMC) for the regression coefficients ($N=100,p=30$) under three allocations of a fixed total posterior sample size ($m\cdot S=4000$) across multiple imputations: m=10,S=400m=10,\,S=400, m=100,S=40m=100,\,S=40, and m=1000,S=4m=1000,\,S=4. Each panel corresponds to one parameter ($b_{\text{Intercept}},b_{V1},\ldots,b_{V10}$). The near-complete overlap of the curves indicates that, in this example, the inferred posterior is largely insensitive to the specific choice of $m$ and $S$ when the total number of draws is held constant.
Table 3: Differences between the individual posterior distributions obtained via MCMC and via the iterative algorithm under the standard prior. Values are the mean absolute differences in the corresponding metric, averaged over 20 datasets, with ($m=100$) imputations per dataset. For each dataset, the two sets of posterior samples are compared using the kernel maximum mean discrepancy (MMD) test at significance level ($\alpha=0.05$). The far-right column reports the percentage of datasets in which the null hypothesis (that the two sample sets are drawn from the same distribution) is rejected.
NN $p$ $\%$ missing rows Mean Standard Deviation $5\%$ Quantile $95\%$ Quantile % MMD rejection
10 2 5 0.0188 0.0177 0.0451 0.0427 0.45
15 0.0241 0.0239 0.0582 0.0578 4.35
30 0.0227 0.0218 0.0532 0.0517 3.10
5 5 0.0432 0.0669 0.1297 0.1213 11.90
15 0.0671 0.0828 0.1520 0.1622 15.30
30 0.0624 0.0719 0.1414 0.1465 14.65
100 10 5 0.0045 0.0036 0.0095 0.0097 0.30
15 0.0069 0.0051 0.0137 0.0141 6.00
30 0.0051 0.0038 0.0106 0.0104 4.45
30 5 0.0077 0.0054 0.0155 0.0157 3.75
15 0.0069 0.0053 0.0145 0.0146 1.90
30 0.0114 0.0000 0.0229 0.0227 10.45
Table 4: Differences between the individual posterior distributions obtained via MCMC and via the iterative algorithm under the horseshoe prior. Values are the mean absolute differences in the corresponding metric, averaged over 20 datasets, with ($m=100$) imputations per dataset. For each dataset, the two sets of posterior samples are compared using the kernel maximum mean discrepancy (MMD) test at significance level ($\alpha=0.05$). The far-right column reports the percentage of datasets in which the null hypothesis (that the two sample sets are drawn from the same distribution) is rejected.
NN $p$ $\%$ missing rows Mean Standard Deviation $5\%$ Quantile $95\%$ Quantile % MMD rejection
10 2 5 0.0156 0.0140 0.0365 0.0331 0.45
15 0.0312 0.0181 0.0482 0.0527 5.20
30 0.0249 0.0167 0.0454 0.0464 6.70
5 5 0.0355 0.0249 0.0651 0.0554 13.35
15 0.0493 0.0333 0.0818 0.0875 23.45
30 0.0575 0.0365 0.0946 0.0900 27.90
100 10 5 0.0044 0.0033 0.0094 0.0095 0.65
15 0.0086 0.0059 0.0171 0.0164 12.10
30 0.0107 0.0066 0.0186 0.0190 38.40
30 5 0.0088 0.0060 0.0175 0.0176 8.45
15 0.0086 0.0048 0.0147 0.0149 36.25
30 0.0001 0.0000 0.0001 0.0001 0.15
Table 5: Differences between the final posterior distributions obtained via MCMC and via the iterative algorithm under the standard prior. Values are the mean absolute differences in the corresponding metric, averaged over 20 datasets. For each dataset, the two sets of posterior samples are compared using the kernel maximum mean discrepancy (MMD) test at significance level ($\alpha=0.05$). The far-right column reports the percentage of datasets in which the null hypothesis (that the two sample sets are drawn from the same distribution) is rejected.
NN $p$ $\%$ missing rows Mean Standard Deviation $5\%$ Quantile $95\%$ Quantile % MMD rejection
10 2 5 0.0141 0.0127 0.0357 0.0299 0
15 0.0155 0.0166 0.0370 0.0394 0
30 0.0127 0.0158 0.0309 0.0335 0
5 5 0.0329 0.0554 0.1033 0.0879 0
15 0.0409 0.0595 0.0868 0.1082 0
30 0.0357 0.0440 0.0763 0.0792 0
100 10 5 0.0029 0.0025 0.0061 0.0066 0
15 0.0034 0.0031 0.0070 0.0076 0
30 0.0027 0.0021 0.0049 0.0048 0
30 5 0.0042 0.0032 0.0083 0.0083 0
15 0.0034 0.0030 0.0066 0.0067 0
30 0.0043 0.0045 0.0088 0.0088 0
Table 6: Differences between the final posterior distributions obtained via MCMC and via the iterative algorithm under the horseshoe prior. Values are the mean absolute differences in the corresponding metric, averaged over 20 datasets. For each dataset, the two sets of posterior samples are compared using the kernel maximum mean discrepancy (MMD) test at significance level ($\alpha=0.05$). The far-right column reports the percentage of datasets in which the null hypothesis (that the two sample sets are drawn from the same distribution) is rejected.
NN $p$ $\%$ missing rows Mean Standard Deviation $5\%$ Quantile $95\%$ Quantile % MMD rejection
10 2 5 0.0109 0.0100 0.0273 0.0223 0
15 0.0240 0.0119 0.0317 0.0360 0
30 0.0183 0.0095 0.0328 0.0274 5
5 5 0.0281 0.0169 0.0466 0.0348 5
15 0.0339 0.0210 0.0490 0.0529 0
30 0.0386 0.0256 0.0557 0.0592 10
100 10 5 0.0028 0.0023 0.0062 0.0062 0
15 0.0039 0.0030 0.0082 0.0078 0
30 0.0026 0.0023 0.0052 0.0047 0
30 5 0.0042 0.0030 0.0085 0.0085 0
15 0.0017 0.0013 0.0033 0.0031 0
30 0.0001 0.0000 0.0001 0.0001 0
@@ Appendix D Additional Results of the Simulation Study for the Surrogate Modeling Problem
Figure 25: Percentage of MCMC runs required to approximate the posteriors of $m=1000$ surrogate parameter draws, for the logistic and PCE surrogate (in rows).